# Vision-driven River Following of UAV via Safe Reinforcement Learning using Semantic Dynamics Model

### v1

@@ 3.4 Constrained Actor Dynamics Estimator Architecture
where $s^{\prime}_{t+h+1}=\mathcal{M}(s_{t+h},a_{t+h})$ is the predicted next state and $a_{t+h}\sim\pi(\cdot|s_{t+h})$. Unlike reward advantages that require a baseline to stabilize updates due to their long-term dependency, cost advantages in safety-constrained RL often prioritize short-horizon, immediate feedback to enforce constraints effectively [73]. Therefore, our formulation avoids bootstrapping by using the predicted cumulative cost itself as the cost advantage estimation. To facilitate exploration while ensuring adherence to safety constraints, the cost advantage function is modulated using a sigmoid transformation. Specifically, a scaling factor $k$ and a cost baseline $c_{b}$ are applied as follows:
@@
This transformation suppresses small predicted costs, preventing overly conservative behavior, while preserving the gradient signal for larger costs, ensuring that the agent remains sensitive to critical safety violations. This non-linear transformation of cost advantage estimation is inspired by softmax temperature heuristics [74] in MCTS for exploration–exploitation trade-off, utility-based risk transforms in risk-sensitive RL [75, 76] and the bounded quadratic structure of Normalized Advantage Functions [77] for stabilizing policy gradients and yielding bounded safety scores for interpretable thresholding. In our experiments, we select $k$ to be 8 and $c_{b}$ to be 0.5 in both environments.
@@ 4.2 Marginal Gain Advantage Estimation Examination
The cost advantage in Equation 3.2 is disabled when calculating the policy gradient (resulting in Equation 22) to isolate the effectiveness of the reward advantage estimator in driving task-related performance. In this comparison, the policy update solely reflects task performance feedback, without being influenced by environmental dangers or cost signals. This allows for a more focused comparison of how different reward advantage estimators contribute to maximizing task-specific rewards. Different advantage estimation methods are compared in their episodic returns during training in the Medium level environment for the same set of 3 seeds, and during testing in the Easy and Hard levels for 30 episodes.
@@ 4.4 Constrained Actor Dynamics Estimator Evaluation
By combining the MGAE method for reward advantage, the SDM for state planning and the cost estimator for cost advantage estimation, we construct the overall Constrained Actor Dynamics Estimator (CADE) architecture, as shown in Figure 3. In CADE, the Lagrangian-based method regulates policy learning by dynamically balancing reward and cost advantages, softly steering the agent toward safer behaviors during training. However, this approach does not enforce strict action constraints during execution. To explore the effectiveness of hard safety interventions, we introduce safety layer baseline: cost-based planning. This safety layer leverages the SDM and the cost estimator to evaluate multiple candidate trajectories before selecting an action, overriding unsafe choices only when necessary.
In cost-based planning, the agent generates a finite set of imagined trajectories, evaluates their cumulative estimated costs, and selects the first action of the trajectory with the lowest cost, enforcing safety constraints during training. The safety layer is activated only after the first third of training to preserve the agent’s exploration ability and are engaged selectively—intervening only if all sampled trajectories starting from the policy-selected action exceed a predefined cost threshold. If no severe failures (high-cost events) are predicted, the safety layer does not override the policy’s decision. The safety layer has not been activated since the beginning of training because the SDM and the cost estimator have not yet converged. The log probability of the selected action is used for importance sampling in policy updates (Equation 3.2). This method provides direct baseline against the Lagrangian-based approach, highlighting the differences between hard safety constraints that intervene at execution time and soft constraints that shape policy updates over time.
@@ 5.3 Constrained Actor Dynamics Estimator
Difficulty Easy Medium Hard
Metrics EpR EpC EpR EpC EpR EpC
MGAE $17.40\pm 4.43$ $0.42\pm 0.46$ $16.30\pm 5.87$ $0.65\pm 0.60$ $15.80\pm 6.12$ $0.75\pm 0.63$
MGAE + Lagrangian $\boldsymbol{19.76\pm 1.12}$ $\boldsymbol{0.33\pm 0.26}$ $\boldsymbol{19.24\pm 2.42}$ $\boldsymbol{0.42\pm 0.30}$ $\boldsymbol{18.86\pm 2.66}$ $\boldsymbol{0.60\pm 0.33}$
MGAE + safety layer $15.20\pm 7.36$ $0.53\pm 0.53$ $14.68\pm 7.44$ $0.72\pm 0.61$ $14.66\pm 6.85$ $0.70\pm 0.54$
@@
Difficulty Easy Medium Hard
Metrics EpR EpC EpR EpC EpR EpC
MGAE $\boldsymbol{85.62\pm 71.48}$ $\boldsymbol{0.95\pm 0.18}$ $\boldsymbol{48.26\pm 38.10}$ $0.98\pm 0.10$ $24.75\pm 26.35$ $\boldsymbol{0.91\pm 0.19}$
MGAE + Lagrangian $36.86\pm 44.66$ $1.00\pm 0.00$ $41.15\pm 33.05$ $0.99\pm 0.07$ $24.14\pm 30.94$ $0.95\pm 0.14$
MGAE + Safety Layer $70.00\pm 58.09$ $0.98\pm 0.09$ $43.07\pm 33.30$ $\boldsymbol{0.96\pm 0.14}$ $\boldsymbol{33.67\pm 35.03}$ $0.93\pm 0.18$
@@ 6 DISCUSSION
Design of Cost Function: In track or river-following tasks, the cost function is designed to be Markovian, relying solely on the current observation, which sufficiently encapsulates immediate environmental risks. This design simplifies the cost estimation process. In scenarios where costs depend on historical context, integrating the cost estimator with a recurrent network could enhance predictive accuracy. Future research could develop adaptive mechanisms that determine the necessity of historical information for cost estimation, adjusting the model architecture accordingly. If using Markovian cost, as we observed the performance discrepancies of safety regulation methods in CliffCircular (Table 4) and SRE (Table 5), correctly relating the visual observation of the agent to the corresponding cost is the key to improving the performance of safety regulation methods during learning.

### v2

@@ 3.4 Constrained Actor Dynamics Estimator Architecture
where horizon actions are induced by the current policy $a_{t+h}\sim\pi_{\theta}(\cdot|s_{t+h})$, $s^{\prime}_{t+h+1}=\mathcal{M}(s_{t+h},a_{t+h})$ is the predicted next state and $\hat{C}(s^{\prime}_{t+h+1})$ is the predicted Markovian cost of this state. Unlike reward advantages that require a baseline to stabilize updates due to their long-term dependency, cost advantages in safety-constrained RL often prioritize short-horizon, immediate feedback to enforce constraints effectively [75]. Therefore, our formulation avoids bootstrapping by using the predicted cumulative cost itself as the cost advantage estimation. We keep this short-horizon form across MDPs for consistency: when costs are terminal-only, the sum reduces to a single step; when costs are dense and state-dependent, using $H>1$ captures near-term hazard accumulation more accurately and provides a better approximation to the cumulative cost-to-go, yielding a more reliable cost-advantage estimate.
To stabilize training and align magnitudes between the estimated reward and cost terms, the cost advantage function is modulated using a sigmoid transformation [^2] We do not claim novelty for this transformation. It is a stabilization choice akin to widely used normalization/clipping heuristics.. Specifically, a scaling factor $k$ and a cost baseline $c_{b}$ are applied as follows:
@@
This transformation suppresses small predicted costs, preventing overly conservative behavior, while preserving the gradient signal for larger costs, ensuring that the agent remains sensitive to critical safety violations. This mirrors common practice in deep RL of normalizing or clipping advantages/rewards to control gradient scales and improve robustness across batches and tasks (e.g., advantage clipping in PPO implementation in Stable-Baselines3 [76]). Besides, keeping cost- and reward-side signals on comparable scales makes the Lagrangian update less sensitive to batch-to-batch variation in raw magnitudes. The constraint mechanism ($\beta$ in Equation 7) remains the primary tool for balancing performance and safety. In our experiments, we select $k$ to be 8 and $c_{b}$ to be 0.5 in both environments.
@@ 4.2 Marginal Gain Advantage Estimation Examination
The cost advantage in Equation 3.2 is disabled when calculating the policy gradient (resulting in Equation 22) to isolate the effectiveness of the reward advantage estimator in driving task-related performance. In this comparison, the policy update solely reflects task performance feedback, without being influenced by environmental dangers or cost signals. This allows for a more focused comparison of how different reward advantage estimators contribute to maximizing task-specific rewards. Window size of MGAE’s baseline (Equation 11) is chosen as $10$ in both environments. Different advantage estimation methods are compared in their episodic returns during training in the Medium level environment for the same set of 3 seeds, and during testing in the Easy and Hard levels for 30 episodes.
@@ 4.4 Constrained Actor Dynamics Estimator Evaluation
By combining the MGAE method for reward advantage, the SDM for state planning and the cost estimator for cost advantage estimation, we construct the overall Constrained Actor Dynamics Estimator (CADE) architecture, as shown in Figure 3. In CADE, the Lagrangian-based method regulates policy learning by dynamically balancing reward and cost advantages, softly steering the agent toward safer behaviors during training. However, this approach does not enforce strict action constraints during execution. To explore the effectiveness of hard safety interventions, we evaluate an alternative safety regulator: cost-based planning, which is a safety layer method (more in Section 2.4 and 3.4). This safety layer is implemented within the CADE framework, leverages the SDM and the cost estimator to evaluate multiple candidate trajectories before selecting an action. It complements the soft Lagrangian training with a hard cost-based action filter that overrides the actor only at selection.
As a safety-regulation strategy evaluated on the same backbone of CADE, for cost-based planning, the agent generates a finite set of imagined trajectories, evaluates their cumulative estimated costs, and selects the first action of the trajectory with the lowest cost, attempting to enforce safety constraints during training. This safety layer does not modify the actor’s training objective, it only overrides actions at selection time using SDM-predicted rollouts and cost estimates, providing a hard alternative to the soft Lagrangian training used in CADE. The safety layer is activated only after the first third of training to preserve the agent’s exploration ability and are engaged selectively—intervening only if all sampled trajectories starting from the policy-selected action exceed a predefined cost threshold. If no severe failures (high-cost events) are predicted, the safety layer does not override the policy’s decision. The safety layer has not been activated since the beginning of training because the SDM and the cost estimator have not yet converged. The log probability of the selected action is used for importance sampling in policy updates (Equation 3.2). This method provides direct comparative against the Lagrangian-based approach, highlighting the differences between hard safety constraints that intervene at execution time and soft constraints that shape policy updates over time.
@@ 5.3 Constrained Actor Dynamics Estimator
MGAE MGAE+Lagrangian MGAE+SafetyLayer
Easy–EpR 17.40 $\pm$ 4.43 19.76 $\pm$ 1.12 15.20 $\pm$ 7.36
Easy–EpC 0.42 $\pm$ 0.46 0.33 $\pm$ 0.26 0.53 $\pm$ 0.53
Medium–EpR 16.30 $\pm$ 5.87 19.24 $\pm$ 2.42 14.68 $\pm$ 7.44
Medium–EpC 0.65 $\pm$ 0.60 0.42 $\pm$ 0.30 0.72 $\pm$ 0.61
Hard–EpR 15.80 $\pm$ 6.12 18.86 $\pm$ 2.66 14.66 $\pm$ 6.85
Hard–EpC 0.75 $\pm$ 0.63 0.60 $\pm$ 0.33 0.70 $\pm$ 0.54
@@
MGAE MGAE+Lagrangian MGAE+SafetyLayer
Easy–EpR $\mathbf{85.62\pm 71.48}$ $36.86\pm 44.66$ $70.00\pm 58.09$
Easy–EpC $\mathbf{0.95\pm 0.18}$ $1.00\pm 0.00$ $0.98\pm 0.09$
Medium–EpR $\mathbf{48.26\pm 38.10}$ $41.15\pm 33.05$ $43.07\pm 33.30$
Medium–EpC $8\pm 0.10 0.99±0$$9\pm 0.07$ $\mathbf{0.96\pm 0.14}$
Hard–EpR $24.75\pm 26.35$ $24.14\pm 30.94$ $\mathbf{33.67\pm 35.03}$
Hard–EpC $\mathbf{0.91\pm 0.19}$ $0.95\pm 0.14$ $0.93\pm 0.18$
@@ 6 DISCUSSION
Design of Cost Function: In track or river following tasks, the cost function is designed to be Markovian, relying solely on the current observation, which sufficiently encapsulates immediate environmental risks. This design simplifies the cost estimation process. In scenarios where costs depend on historical context, integrating the cost estimator with a recurrent network could enhance predictive accuracy. Future research could develop adaptive mechanisms that determine the necessity of historical information for cost estimation, adjusting the model architecture accordingly. If using Markovian cost, as we observed the performance discrepancies of safety regulation methods in CliffCircular (Table 4) and SRE (Table 5), correctly relating the visual observation of the agent to the corresponding cost is the key to improving the performance of safety regulation methods during learning.